# The Smarr formula within the Geroch-Held-Penrose formalism

### v1

@@ Abstract
The connection between classical thermodynamics and black hole horizons is a fundamental topic in gravitational physics, offering a potential pathway to understanding quantum aspects of gravity. However, while black hole mechanics exhibits well-known thermodynamic parallels, a rigorous geometric interpretation of thermodynamic variables directly from the field equations warrants further research. In this manuscript, we present a thermodynamic formulation of the field equations through the decomposition of the Riemann tensor, employing the Geroch-Held-Penrose (GHP) formalism, to clarify a strong correspondence between black hole thermodynamic variables and geometrical quantities derived from horizon geometry. Our analysis reveals an intrinsic connection between the Penrose and Rindler $K$-curvature and the Smarr relation, motivating a revised definition of both trapping gravity and black hole internal energy. Additionally, we derive through this GHP formalism the Smarr formula for the Reissner-Nördstrom black hole cointained in an AdS spacetime and we explore the implications of this relationship for black holes with exotic topologies and in the context of extended theories, exemplified by $f(R)$ gravity. These findings suggest a deeper geometrical basis for black hole thermodynamics, potentially advancing our understanding of gravitational energy, horizon entropy, and their significance within quantum gravity frameworks.
@@ I Introduction
After Hawking’s area theorem for black holes [1], which states that black hole surface cannot decrease [^1] This initial proof assumed asymptotic flatness, the weak energy condition and cosmic censorship. These hypothesis can be restrictive when considering aspects of quantum gravity. Some generalizations of this law have been developed [2, 3]. as a result of any classical process, a new topic in gravitation was developed which tries to study the connections between thermodynamics (possibly from a statistical mechanics perspective) and black hole horizons.
At the same time, Christodoulou showed that black holes possess an irreducible mass which is strongly related to the black hole surface area [4, 5]. This includes even charged Kerr black holes in the context of General Relativity (GR). Building on this, the Penrose inequality conjectures a more general geometric bound: the total ADM mass of a spacetime containing a black hole cannot be less than the irreducible mass associated with the event horizon’s area. Actually, generalized versions of this inequality are still being developed [6].
The black hole area theorem, the existence of an irreducible mass and their similarity with the fundamental notion of classical entropy motivated Bekenstein to obtain an expression for black hole entropy proportional to the area of the horizon [7]. More generally, the black hole entropy was found later to be a Noether charge of the Hamiltonian due to the diffeomorphism invariance property of any [^2] We are assuming that the general covariance principle is necessary to formulate any suitable theory of gravity. theory of gravity [8]. This result extended the study of black hole thermodynamics to more general theories of gravity [9, 10]. In particular, this conserved charge coincides with the Bekenstein-Hawking entropy in GR. Furthermore, quantum corrections to entropy have been studied in different contexts to give a physical interpretation to the entropy of black holes [11].
@@
The entropy associated with an event horizon strongly suggests the existence of its thermodynamic conjugated variable analog to the temperature. Based on this and considering quantum mechanical effects occurring at the horizon, Hawking discovered that black holes emit thermal radiation [12]. More generally, a temperature can be associated to any bifurcation horizon [13] or even to a generic null surface [14].
Black hole thermodynamics, which emerged from these initial results, establishes formal analogies between the behavior of black holes and the fundamental laws of thermodynamics. The pioneering work of Bekenstein and Hawking led to the reformulation of the four laws of black hole mechanics [15] in terms of the classical laws of thermodynamics.
@@
The study of black hole thermodynamics has also motivated a broader look at the thermodynamic properties of horizons in non-black hole contexts, such as in accelerating or cosmological horizons [16, 17], further suggesting that thermodynamic-like descriptions may apply to a wide range of horizon-like structures in gravitational settings.
@@
Within this context, Hayward gave a general quasi-local definition of a black hole in terms of future outer trapping horizons [18]. These trapping horizons are hypersurfaces constructed by a specific foliation of future outer marginal surfaces, which are Riemannian 2-surfaces where a null expansion vanishes (these marginal surfaces are the 2-dimensional horizons usually seen in pictures or movies). Moreover, a generalization of the surface gravity, called trapping gravity, and the related black hole dynamic laws are given, which were also translated into GHP formalism, allowing a more geometric interpretation [19].
A key limitation is that Hayward’s definition for trapping gravity (see Eq. (7) in [19]), does not coincide with the surface gravity even in the context of static and spherically symmetric spacetimes, e.g. the Reissner-Nördstrom black hole [20, 21]. A revised definition is proposed in this manuscript based in the first law of black hole thermodynamics.
Finally, let us comment that connections between thermodynamics and gravity have been explored at a fundamental level. Jacobson derived Einstein equations as a consequence of the first law of thermodynamics [22], highlighting their intrinsic relationship. Conversely, Padmanabhan demonstrated that the first law of thermodynamics can also be derived from Einstein’s equations in certain simplified scenarios [23], underlying the deep connection between gravity and thermodynamics.
This manuscript is organized as follows: Sec. II provides a brief overview of Padmanabhan’s results connecting the Einstein field equations to the first law of black hole thermodynamics in static and spherically symmetric spacetimes. In Sec. III we present the GHP formulation of the field equations, analyzing the nontrivial equations depending on the underlying geometric structure of the spacetime and their relation with geometrical quantities. In Sec. IV, we present the key findings of this work, exploring thermodynamics within the GHP formalism for static spacetimes foliated by compact orientable 2-surfaces of constant Gaussian curvature. Eventually, in Sec. V we present some final remarks and interesting directions for future research.
@@ II Padmanabhan’s view of Einstein field equations as a thermodynamic relation
Initially, the laws of black hole thermodynamics were constructed without mentioning the Einstein field equations. Surprisingly, Padmanabhan showed that the first law of black hole thermodynamics is also encoded in the field equations [23]. This was done first in the context of GR for static and spherically symmetric spacetimes with line element
@@
where the pressure is defined as $P=-T_{r}^{r}$. Moreover, this identification is uniquely fixed [10].
The corresponding Smarr relation can be obtained by using a scaling argument and Euler’s theorem [24] to integrate Eq. (5). Note that $E\sim r$, $S\sim r^{2}$ and $V\sim r^{3}$. Thus, the Smarr formula is given by
@@
Note that this relation is not exactly the same as the Smarr mass formula for charged Kerr black holes [25]. In fact, Eq. (5) and its corresponding integrated form shown in Eq. (6) generalize the Smarr formula for any static and spherically symmetric black hole in GR.
For the vacuum case and without cosmological constant, $E$ coincides with the mass of the black hole and Eq. (6) reduces to $M=2TS$. Alternatively, $E$ does not coincide with the mass of the black hole when considering a cosmological constant [26]. Actually, when the pressure term is due to the presence of a negative cosmological constant, $P=-\lambda/8\pi$, the variable $E$ denotes the total internal energy of the black hole and its mass may be identified with the black hole chemical enthalpy $M=E+PV$ [27]. In such a case, the Smarr relation in Eq. (6) takes the form
@@
Finally, let us point out that this derivation was extended to any static spacetime for Lanczos-Lovelock gravity [28] and for stationary axisymmetric horizons [29].
@@ III GHP field equations
Let us assume that the spacetime $(M,g)$ is defined as a $\mathcal{C}^{k}$-manifold and the metric tensor is of type $\mathcal{C}^{k-1}$. This assumption on differentiability allows the field equations of a extended theory of gravity of order $k-1$ to be well-defined everywhere on $M$. Through all the manuscript we use the signature (1,3) for the metric and the Penrose and Rindler’s sign convention for the Riemann curvature tensor [30].
@@
where $\Delta t_{\mu\nu}$ are the components of some effective stress-energy tensor which does not include the Ricci tensor. Many general Lagrangians allow the field equations to take the form of Eq. (8). Some examples include GR, $f(R)$ gravity, scalar-tensor theories, Lovelock theories, etc. To the best of our knowledge, there has been no comprehensive analysis investigating which classes of Lagrangian allow this specific formulation of the field equations. A systematic study in this direction might not only broaden our understanding of existing theoretical frameworks but also inspire the development of new theories of gravity.
Within the spin coefficient formalism, and also in the GHP form, the so called Ricci scalars, $\Phi_{ij}$ with $i,j\in\{0,1,2\}$, are defined as the contractions of the Ricci tensor with a null tetrad. These are the components of the Ricci tensor in the new basis [31, 32]. Thus, the corresponding field equations, Eqs. (8), in this new basis can be written as $\Phi_{ij}=\Phi_{ij}^{ph}$ and $\Lambda=\Lambda^{ph}$, where $\Phi_{ij}^{ph}$ define the physical contractions of the effective stress-energy tensor in an analogous way as the Ricci scalars and $\Lambda^{ph}=\Delta t^{\mu}_{\mu}/24$, where $\Delta t^{\mu}_{\mu}$ is the trace of the effective stress-energy tensor.
At this point, the field equations in the new basis can be rewritten in terms of geometrical quantities by using the following GHP equations and the definition of the Penrose-Rindler $K$-curvature (see Eq. (4.14.20) in [30]). That is, the field equations are
@@
plus the respective field equations obtained from applying prime (), complex conjugate (c.c.) or Sachs ($\ast$) operators recursively as the diagrams in Fig. 1 shows,
@@
where $(i,j)$ denotes $\Phi_{ij}$. For instance, $(1,1)_{\pm 3\Lambda}$ means $\Phi_{11}\pm 3\Lambda$.
@@
$\mbox{}\rho^{\prime}-\eth\tau^{\prime}$ $=$ $\rho\overline{\rho}^{\prime}+\sigma\sigma^{\prime}-\tau^{\prime}\overline{\tau}^{\prime}-\kappa\kappa^{\prime}-(\Psi_{2}+2\Lambda),$ (13)
Based on Hayward’s definition of trapping gravity [19],
@@
is the Gaussian curvature of the considered spacelike 2-surfaces (see Proposition 4.14.21 in [30] and App. A). That is, $k_{g}$ is the Gaussian curvature of the future outer marginal surface when evaluating Eq. (15) at the horizon [^10] From now on, by horizon we are referring to the a future outer marginal surfaces which foliates the trapping horizon in the Hayward’s definition of black hole [19]. Do not confuse with the global definition of event horizon. Essentially, this is a cross section of the trapping horizon. This terminology is widely used in the literature.. Using this equation, Eqs. (15) and (16) can be rewritten as
@@
This allows a geometric interpretation of the definition of black hole horizon, characterized by the equilibrium between material, rotation and shear energy densities on the right hand side and a surface tension term on the left hand side. This relation suggest that $\tilde{k}_{g}$ has an interpretation of energy density and $K-K^{*}$ has dynamical meaning, which was pointed out by Hayward [19].

### v2

@@ Abstract
The connection between classical thermodynamics and black hole horizons is a fundamental topic in gravitational physics, offering a potential pathway to understanding quantum aspects of gravity. However, while black hole mechanics exhibits well-known thermodynamic parallels, a rigorous geometric interpretation of thermodynamic variables directly from the field equations warrants further research. In this manuscript, we present a thermodynamic formulation of the field equations through the decomposition of the Riemann tensor, employing the Geroch-Held-Penrose (GHP) formalism, to clarify a strong correspondence between black hole thermodynamic variables and geometrical quantities derived from horizon geometry. Our analysis reveals an intrinsic connection between the Penrose and Rindler $K$-curvature and the Smarr relation, motivating a revised definition of both trapping gravity and black hole internal energy. Additionally, we derive through this GHP formalism the Smarr formula for the Reissner-Nördstrom black hole cointained in an AdS spacetime and we explore the implications of this relationship for black holes with exotic topologies and in the context of extended theories, exemplified by $f(R)$ gravity. We also briefly analyze this relation in more general configurations, such as Taub-NUT spacetimes. These findings suggest a deeper geometrical basis for black hole thermodynamics, potentially advancing our understanding of gravitational energy, horizon entropy, and their significance within quantum gravity frameworks.
@@ I Introduction
After Hawking’s area theorem for black holes Hawking1972 , which states that black hole surface cannot decrease [^1] This initial proof assumed asymptotic flatness, the weak energy condition and cosmic censorship. These hypothesis can be restrictive when considering aspects of quantum gravity. Some generalizations of this law have been developed Woolgar1999 ; Kontou2024 . as a result of any classical process, a new topic in gravitation was developed which tries to study the connections between thermodynamics (possibly from a statistical mechanics perspective) and black hole horizons.
At the same time, Christodoulou showed that black holes possess an irreducible mass which is strongly related to the black hole surface area Christodoulou1970 ; Christodoulou1971 . This includes even charged Kerr black holes in the context of General Relativity (GR). Building on this, the Penrose inequality conjectures a more general geometric bound: the total ADM mass of a spacetime containing a black hole cannot be less than the irreducible mass associated with the event horizon’s area. Actually, generalized versions of this inequality are still being developed Harvie2024 .
The black hole area theorem, the existence of an irreducible mass and their similarity with the fundamental notion of classical entropy motivated Bekenstein to obtain an expression for black hole entropy proportional to the area of the horizon Bekenstein1973 . More generally, the black hole entropy was found later to be a Noether charge of the Hamiltonian due to the diffeomorphism invariance property of any [^2] We are assuming that the general covariance principle is necessary to formulate any suitable theory of gravity. theory of gravity Wald1993 . This result extended the study of black hole thermodynamics to more general theories of gravity Faraoni2010 ; Padmanabhan2010 . In particular, the thermodynamics of $f(R)$ gravity has been extensively studied, including various formulations of this theory Bamba2010 ; Peralta2020 . This conserved charge coincides with the Bekenstein-Hawking entropy in GR. Furthermore, quantum corrections to entropy have been studied in different contexts to give a physical interpretation to the entropy of black holes Susskind1993 .
@@
The entropy associated with an event horizon strongly suggests the existence of its thermodynamic conjugated variable analog to the temperature. Based on this and considering quantum mechanical effects occurring at the horizon, Hawking discovered that black holes emit thermal radiation Hawking1975 . More generally, a temperature can be associated to any bifurcation horizon Lee1986 or even to a generic null surface Dalui2021 .
Black hole thermodynamics, which emerged from these initial results, establishes formal analogies between the behavior of black holes and the fundamental laws of thermodynamics. The pioneering work of Bekenstein and Hawking led to the reformulation of the four laws of black hole mechanics Bardeen1973 in terms of the classical laws of thermodynamics.
@@
The study of black hole thermodynamics has also motivated a broader look at the thermodynamic properties of horizons in non-black hole contexts, such as in accelerating or cosmological horizons Appels2016 ; Gibbons1977b , further suggesting that thermodynamic-like descriptions may apply to a wide range of horizon-like structures in gravitational settings.
@@
Within this context, Hayward gave a general quasi-local definition of a black hole in terms of future outer trapping horizons Hayward1994d . These trapping horizons are hypersurfaces constructed by a specific foliation of future outer marginal surfaces, which are Riemannian 2-surfaces where a null expansion vanishes (these marginal surfaces are the 2-dimensional horizons usually seen in pictures or movies). Moreover, a generalization of the surface gravity, called trapping gravity, and the related black hole dynamic laws are given, which were also translated into GHP formalism, allowing a more geometric interpretation Hayward1994 .
A key limitation is that Hayward’s definition for trapping gravity (see Eq. (7) in Hayward1994 ), does not coincide with the surface gravity even in the context of static and spherically symmetric spacetimes, e.g. the Reissner-Nördstrom black hole Fodor1996 ; Nielsen2008 . A revised definition is proposed in this manuscript based in the first law of black hole thermodynamics.
Finally, let us comment that connections between thermodynamics and gravity have been explored at a fundamental level. Jacobson derived Einstein equations as a consequence of the first law of thermodynamics Jacobson1995 , highlighting their intrinsic relationship. In addition, the field equations of $f(R)$ gravity were derived in Eling2006 , where the Clausius relation was extended to non-equilibrium thermodynamics within this framework.
Conversely, Padmanabhan demonstrated that the first law of thermodynamics can also be derived from Einstein’s equations in certain simplified scenarios Padmanabhan2002 , underlying the deep connection between gravity and thermodynamics.
The study of black hole thermodynamics has revealed deep connections between gravity, thermodynamics, and quantum theory. However, a key challenge arises from the global nature of event horizons, which limits their physical applicability in realistic cosmological settings. Quasi-local approaches, such as Hayward’s trapping horizons, offer a more general framework, but existing definitions of trapping gravity fail to recover the expected thermodynamic properties in well-known cases. Based on Padmanabhan’s approach, we propose a revised definition of trapping gravity inspired by the first law of black hole thermodynamics, aiming to provide a more consistent description within the Geroch-Held-Penrose (GHP) formalism. By formulating black hole thermodynamics in a purely geometric language, we hope to deepen our understanding of the fundamental relationship between gravity and thermodynamics and explore its broader implications.
This manuscript is organized as follows: Sec. II provides a brief overview of Padmanabhan’s results connecting the Einstein field equations to the first law of black hole thermodynamics in static and spherically symmetric spacetimes. In Sec. III we present the GHP formulation of the field equations, analyzing the nontrivial equations depending on the underlying geometric structure of the spacetime and their relation with geometrical quantities. In Sec. IV, we present the key findings of this work, exploring thermodynamics within the GHP formalism for static spacetimes foliated by compact orientable 2-surfaces of constant Gaussian curvature. Eventually, in Sec. VI we present some final remarks and interesting directions for future research.
@@ II Padmanabhan’s view of Einstein field equations as a thermodynamic relation
Initially, the laws of black hole thermodynamics were constructed without mentioning the Einstein field equations. Surprisingly, Padmanabhan showed that the first law of black hole thermodynamics is also encoded in the field equations Padmanabhan2002 . This was done first in the context of GR for static and spherically symmetric spacetimes with line element
@@
where the pressure is defined as $P=-T_{r}^{r}$. Moreover, this identification is uniquely fixed Padmanabhan2010 .
The corresponding Smarr relation can be obtained by using a scaling argument and Euler’s theorem Kastor2009 to integrate Eq. (5). Note that $E\sim r$, $S\sim r^{2}$ and $V\sim r^{3}$. Thus, the Smarr formula is given by
@@
Note that this relation is not exactly the same as the Smarr mass formula for charged Kerr black holes Smarr1973 . In fact, Eq. (5) and its corresponding integrated form shown in Eq. (6) generalize the Smarr formula for any static and spherically symmetric black hole in GR.
For the vacuum case and without cosmological constant, $E$ coincides with the mass of the black hole and Eq. (6) reduces to $M=2TS$. Alternatively, $E$ does not coincide with the mass of the black hole when considering a cosmological constant Kubizk2017 . Actually, when the pressure term is due to the presence of a negative cosmological constant, $P=-\lambda/8\pi$, the variable $E$ denotes the total internal energy of the black hole and its mass may be identified with the black hole chemical enthalpy $M=E+PV$ Cveti2011 . In such a case, the Smarr relation in Eq. (6) takes the form
@@
Finally, let us point out that this derivation was extended to any static spacetime for Lanczos-Lovelock gravity Kothawala2009 and for stationary axisymmetric horizons Kothawala2007 .
@@ III GHP field equations
Let us assume that the spacetime $(M,g)$ is defined as a $\mathcal{C}^{k}$-manifold and the metric tensor is of type $\mathcal{C}^{k-1}$. This assumption on differentiability allows the field equations of a extended theory of gravity of order $k-1$ to be well-defined everywhere on $M$. Through all the manuscript we use the signature (1,3) for the metric and the Penrose and Rindler’s sign convention for the Riemann curvature tensor Penrose1984 .
@@
where $\Delta t_{\mu\nu}$ are the components of some effective stress-energy tensor which does not include the Ricci tensor. Many general Lagrangians allow the field equations to take the form of Eq. (8). Some examples include GR, $f(R)$ gravity, scalar-tensor theories, etc. For the aforementioned examples we have: (i) within GR, the effective stress energy tensor is given by $\Delta t_{\mu\nu}^{(GR)}=T_{\mu\nu}^{(m)}-\frac{1}{2}T^{(m)}g_{\mu\nu}+\lambda g_{\mu\nu}$, where $T_{\mu\nu}^{(m)}$ is the stress-energy tensor associated with the material Lagrangian and $\lambda$ is the cosmological constant; (ii) the effective stress energy tensor in $f(R)$ gravity is given by
where $F(R)\coloneqq df(R)/dR$; and (iii) for scalar-tensor theories it holds that
$\Delta t_{\mu\nu}^{(\phi)}=\dfrac{1}{\phi}T_{\mu\nu}^{(m)}-\dfrac{1}{2}Rg_{\mu\nu}+\dfrac{1}{\phi}\left[\nabla_{\mu}\nabla_{\nu}-g_{\mu\nu}\square\right]\phi+\dfrac{\omega(\phi)}{\phi^{2}}\left(\partial_{\mu}\phi\partial_{\nu}\phi+\dfrac{1}{2}g_{\mu\nu}\partial_{\sigma}\phi\partial^{\sigma}\phi\right),$ (10)
where $\phi$ is an auxiliar (scalar) field which must be a solution of the corresponding scalar-field equation Quiros2019 .
To the best of our knowledge, there has been no comprehensive analysis investigating which classes of Lagrangian allow this specific formulation of the field equations. A systematic study in this direction might not only broaden our understanding of existing theoretical frameworks but also inspire the development of new theories of gravity.
In alignment with the spin coefficient and GHP formalisms, we construct a null tetrad consisting of two null vectors, $\mathbf{l}$ and $\mathbf{n}$. For the cases examined in this manuscript, these vectors can be directly aligned with the principal null directions NewmanPenrose . The tetrad is completed by introducing a complex null vector $\mathbf{m}$ and its conjugate $\overline{\mathbf{m}}$, formed by combining two real, orthonormal spacelike vectors.
Within the spin coefficient formalism, and also in the GHP form, the so called Ricci scalars, $\Phi_{ij}$ with $i,j\in\{0,1,2\}$, are defined as the contractions of the Ricci tensor with the null tetrad. These are the components of the Ricci tensor in the new basis NewmanPenrose ; GHPBook . Thus, the corresponding field equations, Eq. (8), in this new basis can be written as $\Phi_{ij}=\Phi_{ij}^{ph}$ and $\Lambda=\Lambda^{ph}$, where $\Phi_{ij}^{ph}$ define the physical contractions of the effective stress-energy tensor in an analogous way as the Ricci scalars and $\Lambda^{ph}=\Delta t/24$, where $\Delta t=\Delta t^{\mu}_{\mu}$ is the trace of the effective stress-energy tensor.
At this point, the field equations in the new basis can be rewritten in terms of geometrical quantities by using the GHP equations. Here, (thorn), $\eth$ (edth) and their primed form denote the spin weighted derivatives, related with the directional derivatives along the null tetrad vectors, respectively, where primes () indicate transformations under the exchange $\mathbf{l}\leftrightarrow\mathbf{n}$ and $\mathbf{m}\leftrightarrow\overline{\mathbf{m}}$.
and the definition of the Penrose-Rindler $K$-curvature (see Eq. (4.14.20) in Penrose1984 ). That is, the field equations are
@@
where overline denote complex conjugation and $\Psi_{2}$ is the Weyl curvature scalar
$\Psi_{2}=C_{\alpha\beta\gamma\delta}l^{\alpha}m^{\beta}\overline{m}^{\gamma}n^{\delta}.$ (14)
Additional field equations are obtained by systematically applying prime (), complex conjugate (c.c.) or Sachs ($\ast$) operators recursively as the diagrams in Fig. 1 shows,
@@
where $(i,j)$ denotes $\Phi_{ij}$. For instance, $(1,1)_{\pm 3\Lambda}$ means $\Phi_{11}\pm 3\Lambda$. Here, the Sachs operator ($\ast$) is defined by the following transformations:
$\mathbf{l}^{\ast}=\mathbf{m},\quad\mathbf{n}^{\ast}=-\overline{\mathbf{m}},\quad\mathbf{m}^{\ast}=-\mathbf{l},\quad\overline{\mathbf{m}}^{\ast}=\mathbf{n}$ (15)
For further details check references Penrose1984 ; Geroch1973 ; GHPBook .
@@
$\mbox{}\rho^{\prime}-\eth\tau^{\prime}$ $=$ $\rho\overline{\rho}^{\prime}+\sigma\sigma^{\prime}-\tau^{\prime}\overline{\tau}^{\prime}-\kappa\kappa^{\prime}-(\Psi_{2}+2\Lambda).$ (17)
Based on Hayward’s definition of trapping gravity Hayward1994 ,
@@
is the Gaussian curvature of the considered spacelike 2-surfaces (see Proposition 4.14.21 in Penrose1984 and App. A). That is, $k_{g}$ is the Gaussian curvature of the future outer marginal surface when evaluating Eq. (19) at the horizon [^10] From now on, by horizon we are referring to the future outer marginal surfaces which foliates the trapping horizon in the Hayward’s definition of black hole Hayward1994 . Do not confuse with the global definition of event horizon. Essentially, this is a cross section of the trapping horizon. This terminology is widely used in the literature.. Using this equation, Eqs. (19) and (20) can be rewritten as
@@
This allows a geometric interpretation of the definition of black hole horizon, characterized by the equilibrium between material, rotation and shear energy densities on the right hand side and a surface tension term on the left hand side. This relation suggest that $\tilde{k}_{g}$ has an interpretation of energy density and $K-K^{*}$ has dynamical meaning, which was pointed out by Hayward Hayward1994 .